# Beyond Multiple Choice: Evaluating Steering Vectors for Adaptive Free-Form Summarization

### v1

@@ 1 Introduction
The use of large pre-trained language models, which have been trained on extensive web-based datasets, has emerged as the preferred method for addressing numerous tasks in natural language processing (NLP) (Devlin et al., 2019; Brown et al., 2020). Consequently, the ability to adapt foundation models to specific tasks and align their outputs with user preferences has become crucial. Previous research on controlling language models can often be classified into three main strategies: prompt engineering (Shin et al., 2020; Lester et al., 2021; Wei et al., 2022), trainable decoding mechanisms (Deng et al., 2020) and fine-tuning according to specific objectives (Ouyang et al., 2022a; Rafailov et al., 2023).
@@
    We release our source code and steering vector training datasets to promote reproducibility and facilitate further research, available at: Anonymized Repository.
@@ 2 Related Work
Generating adaptive summaries tailored to user preferences typically involves fine-tuning existing foundation models, modifying model architectures, or employing specialized training procedures (Urlana et al., 2024; Bahrainian et al., 2024; Zhang et al., 2025). For instance, (Bahrainian et al., 2021) introduces an abstractive summarization model that enables topic-level customization through a novel ’topical attention’ mechanism. Similarly, (Blinova et al., 2023) proposes a two-stage model for document-level text simplification that first summarizes and then further simplifies content using transformers, enhanced by keyword prompts and an embedding similarity loss. (Bahrainian et al., 2023) use a Transformer-based architecture for controllable topic-focused summarization, which modifies the cross-attention mechanism for guiding the topical focus.
@@
Despite their appeal as lightweight control methods, activation steering methods face significant challenges (Braun et al., 2024). Recent studies highlight issues with reliability and generalization, noting high variance across inputs and instances where steering produces the opposite of the intended effect (Tan et al., 2024; Brumley et al., 2024; Braun et al., 2025). Furthermore, steering vectors are often evaluated in constrained settings, like multiple-choice questions, rather than more challenging free-form generation tasks (Pres et al., 2024; Braun et al., 2024)
@@ 5 Discussion
This study evaluates the effectiveness of CAA steering vectors for controlling relevant text properties during free-form abstractive summarization. Our findings demonstrate that steering effectively controls topical focus, sentiment, and readability, but this control inherently involves an efficacy-quality trade-off: higher steering strengths achieve greater control at the cost of significant degradation in both intrinsic and extrinsic summary quality.
@@
Compared to steering, prompt engineering offered weaker control but substantially better preservation of text quality. This makes prompting a viable alternative when quality is paramount and moderate control suffices. Combining steering vectors with prompting emerged as the most promising strategy, yielding the strongest control, often already with moderate steering strengths. This hybrid approach achieved the most favorable efficacy-quality balance, though large steering strengths still degrade text quality substantially.
@@ 5.1 Limitations
Our conclusions are shaped and limited by our key methodological choices. We only use CAA steering vectors and our findings may not generalize across all steering methods. Similarly, the results are specific to the summarization task on the NEWTS dataset and the Llama model family. Performance on other tasks, datasets, or model architectures could differ. Furthermore, the automated metrics used for evaluation, while standard, have inherent limitations in fully capturing nuanced human judgments. Broader research is therefore necessary to further validate the effectiveness of steering methods for free-form generation tasks.
@@ 5.2 Future Work
The observed trade-off between control efficacy and text quality degradation motivates methods that find an optimal control-quality trade-off. Developing a decision mechanism to dynamically adjust the steering strength $\lambda$ could be promising. For instance, one could project the incoming activation onto a linear classifier trained on the steering vector training data and only apply the steering vector with the strength needed to shift the activation to the desired side of the decision boundary. Such an approach could potentially minimize text quality degradation while maintaining strong control over text properties by applying steering only when necessary and only as much as necessary.
Another important area for future exploration is the application of steering vectors in multiple-attribute controllable summarization. This would involve developing and applying methods to steer multiple text properties simultaneously. Such an approach could present new challenges related to vector composition, potential interference between steering directions, and managing cumulative impacts on text quality.
@@ C.11 Impact across Model Sizes
The impact of prompting across model scales can be found in Aooendix C.4.

### v2

@@ 1 Introduction
Large pre-trained language models have emerged as the preferred method for addressing numerous natural language processing (NLP) tasks (Devlin et al., 2019; Brown et al., 2020). Consequently, the ability to adapt foundation models to specific tasks and align their outputs with user preferences is crucial. Previous research on controlling language models can often be classified into three main strategies: prompt engineering (Shin et al., 2020; Lester et al., 2021; Wei et al., 2022), trainable decoding mechanisms (Deng et al., 2020) and fine-tuning according to specific objectives (Ouyang et al., 2022a; Rafailov et al., 2023).
@@
    We release our source code and steering vector training datasets to promote reproducibility and facilitate further research, available at: GitHub Repository.
@@ 2 Related Work
Generating adaptive summaries tailored to user preferences typically involves fine-tuning existing foundation models, modifying model architectures, or employing specialized training procedures (Urlana et al., 2024; Bahrainian et al., 2024; Zhang et al., 2025; Braun et al., 2025b). For instance, Bahrainian et al. (2021) introduces an abstractive summarization model that enables topic-level customization through a novel ’topical attention’ mechanism. Similarly, Blinova et al. (2023) proposes a two-stage model for document-level text simplification that first summarizes and then further simplifies content using transformers, enhanced by keyword prompts and an embedding similarity loss. Bahrainian et al. (2023) use a Transformer-based architecture for controllable topic-focused summarization, which modifies the cross-attention mechanism for guiding the topical focus.
@@
Despite their appeal as lightweight control methods, activation steering methods face significant challenges (Braun et al., 2024). Recent studies highlight issues with reliability and generalization, noting high variance across inputs and instances where steering produces the opposite of the intended effect (Tan et al., 2024; Brumley et al., 2024; Braun et al., 2025a). Furthermore, steering vectors are often evaluated in constrained settings, like multiple-choice questions, rather than more challenging free-form generation tasks (Pres et al., 2024; Braun et al., 2024)
@@ 5 Discussion
This study evaluates the effectiveness of CAA steering vectors for controlling relevant text properties during free-form abstractive summarization. Our findings demonstrate that steering effectively controls topical focus, sentiment (Bahrainian & Dengel, 2015), and readability, but this control inherently involves an efficacy-quality trade-off: higher steering strengths achieve greater control at the cost of significant degradation in both intrinsic and extrinsic summary quality.
@@
Compared to steering, prompt engineering offered weaker control but substantially better preservation of text quality. This makes prompting a viable alternative when quality is paramount and moderate control suffices. Combining steering vectors with prompting emerged as the most promising strategy, yielding the strongest control, often already with moderate steering strengths. This hybrid approach achieved the most favorable efficacy-quality trade-off, though large steering strengths still degrade text quality substantially.
@@ 5.1 Limitations
Our conclusions are shaped and limited by our key methodological choices. We only use CAA steering vectors and our findings may not generalize across all steering methods. Similarly, the results are specific to the summarization task on the NEWTS dataset and the Llama model family. Performance in other tasks, data sets, or model architectures could differ. Furthermore, the automated metrics used for evaluation, while standard, have inherent limitations in fully capturing nuanced human judgments. Broader research is therefore necessary to further validate the effectiveness of steering methods for free-form generation tasks.
@@ 5.2 Future Work
The observed trade-off between control efficacy and text quality degradation motivates methods that find an optimal trade-off between control quality and control. Developing a decision mechanism to dynamically adjust the steering strength $\lambda$ could be promising. For example, one could project the incoming activation onto a linear classifier trained on the steering vector training data and only apply the steering vector with the strength needed to shift the activation to the desired side of the decision boundary. Such an approach could potentially minimize text quality degradation while maintaining strong control over text properties by applying steering only when necessary and only as much as necessary.
Another important area for future exploration is the application of steering vectors in multiple-attribute controllable summarization. This would involve developing and applying methods to steer multiple text properties simultaneously. This approach could present new challenges related to vector composition, possible interference between steering directions, and managing cumulative impacts on text quality.
@@ 5.3 Conclusion
Acknowledgements
We thank the anonymous reviewers for their constructive feedback which helped to improve the manuscript. This research utilized compute resources at the Tübingen Machine Learning Cloud, DFG FKZ INST 37/1057-1 FUGG.
@@ C.11 Impact across Model Sizes
The impact of prompting across model scales can be found in Appendix C.4.